OpenStudy (compassionate):

Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. Use the function: F(x) = 12/(4x+2). f(x) = 12/(4x+2) y = 12/(4x+2) 4x+2=12/y 4x = (12/y) - 2 x = 3/y -1/2 Now, we have solved for x, we would plug that into f(h(x)), right? Or no? I'm really confused guize.

OpenStudy (shamil98):

yah

OpenStudy (shamil98):

totes mcgoats

OpenStudy (compassionate):

h(x) = 3/y -1/2, would that be correct...?

OpenStudy (shamil98):

yes, you're so smart!

OpenStudy (compassionate):

I now I have f(h(x))... Are you being sarcastic?

OpenStudy (shamil98):

Not at all, i truly believe you've done this well, well done :)

OpenStudy (the_fizicx99):

I think Comp's onto something :3

OpenStudy (compassionate):

I have the feeling I'm being trolled because this is such basic math but my mind cannot comprehend this stuff.

OpenStudy (compassionate):

What would be my next then then...?

OpenStudy (shamil98):

Not at all, now just find h(f(x)) :)

OpenStudy (compassionate):

That's what I'm trying to find and don't know how...

OpenStudy (compassionate):

I just solved for x, right? In that equation... so would that be h(x) then, right? But how? And wouldn't doing that just give me the same thing?

OpenStudy (shamil98):

Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. Use the function: F(x) = 12/(4x+2). You're supposed to create a new function, not solve for x. I've mislead you mostly because I feel that you are trolling people with basic math questions. Anyways, let's say h(x) = 5x + 3 "assign any number to x" let x = 1 so we have h(1) = 5x + 3 replace the x. h(1) = 5(1) + 3 = 8 Now find out what f(h(x))

OpenStudy (jdoe0001):

\(\bf f(x)=y=\cfrac{12}{4x+2}\qquad inverse\implies x=\cfrac{12}{4y+2}\\ \quad \\\to y=\cfrac{3}{x}-\cfrac{1}{2}\leftarrow f^{-1}(x) \\ \quad \\ thus\qquad f(\quad f^{-1}(x)\quad )=x\quad and\quad f^{-1}(\quad f(x)\quad )=x\)

OpenStudy (compassionate):

I'm not in the slightest trolling. I don't understand basic math.

OpenStudy (jdoe0001):

you found the inverse of your function... now just finding the say f(h(x)) will give "x", since h(x) and f(x) are inverse of each other

OpenStudy (compassionate):

So to find f(h(x)), We could say: 12/(4x+2 = 5x + 3? Or would be add them or?

OpenStudy (compassionate):

The inverse? I'd have to swap x and y for the inverse, right?

OpenStudy (jdoe0001):

ohh ahemm nope... notice shamil98's lines above so keep in mind that \(\bf f(x)=y=\cfrac{12}{4x+2}\qquad inverse\implies x=\cfrac{12}{4y+2} \\ \quad \\ \to y=\cfrac{3}{x}-\cfrac{1}{2}\leftarrow h(x) \\ \quad \\ thus \quad f(x)=\cfrac{12}{4x+2}\qquad {\color{brown}{ h(x)}}=\cfrac{3}{x}-\cfrac{1}{2} \\ \quad \\ f(\quad h(x)\quad )=\cfrac{12}{4{\color{brown}{ h(x)}}+2}\to \cfrac{12}{4\left({\color{brown}{ \frac{3}{x}-\frac{1}{2}}}\right)+2}\)

OpenStudy (compassionate):

@jdoe0001 , I understand what you're saying. And I see I didn't create a a new function, and Shamil told me to solve: f(h(x)), right? But Shamil, you also put "1" in your function ( h(1) = 5x + 3 ) and solved it. Would I need to take F(x) = 12/(4x+2) and add "1" as well, then just add or multiply them like normal?

OpenStudy (compassionate):

You lost me at h(x) = 3/x - 1/2. How does that happen? Where did h(x) come from? Wasn't it f(x) two seconds ago?

OpenStudy (shamil98):

Jdoe is better versed in this area of basic math, my brain is starting to hurt (thinking :/) I shall leave this to him to guide you, Compassionate. All the best in your journey of learning basic mathematics.

OpenStudy (jdoe0001):

hehhe

OpenStudy (compassionate):

It's funny because I can integrate and do derivatives but this stuff throws me off.

OpenStudy (jdoe0001):

well... I got the inverse... by swapping the variables first and then solving for "y"

OpenStudy (jdoe0001):

which is what I thought you did lemme redo the inverse part, so you see it

OpenStudy (compassionate):

Okay, wait, both you guys confused me. Lets start from the top. "Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. Use the function: F(x) = 12/(4x+2)" Okay, so lets say \[f(x) = \frac{ 12 }{ 4x+2 }\]\[h(x) = 3x + 5\] Now I have both my functions, and I have to decide whether f(h(x)) and h(f(x)) will result in the same number. So, how would I put these into f(h(x)) and h(f(x))

OpenStudy (jdoe0001):

hmm

OpenStudy (compassionate):

Remember, h(x) is the function I had created just now like the question is asking.

OpenStudy (jdoe0001):

so f(x) is given and h(x) is just made up, ok

OpenStudy (compassionate):

Yes. It asked me to make up something for h(x) then gave me the two function-thingies above.

OpenStudy (jdoe0001):

well... if ahemm the issue is that if you plug any number into f( h(x) ) and the same number into h( f(x) ) you will NOT get the same value

OpenStudy (compassionate):

Can you explain why... :)

OpenStudy (jdoe0001):

keeping in mind that \(\bf f(\quad h(x)\quad )=\cfrac{12}{4({\color{brown}{ 3x+5}})+2}\qquad h(\quad f(x)\quad )=3\left({\color{brown}{ \cfrac{12}{4x+2}}}\right)+5\) easy example is... as shamil98 said.. set x= 1 and you'll see what'd get

OpenStudy (jdoe0001):

f( h(x) ) = x and h( f(x) ) = x ONLY if they're inverse of each other

OpenStudy (compassionate):

Hmm. Yes, I see. Another question. What happened to 4x?

OpenStudy (compassionate):

In your first step we have 12/4(3x + 5) + 2, wasn't it 4x?

OpenStudy (jdoe0001):

notice the argument for f(x) is NOT "x" is h(x), notice the inserted "brown" h(x) as the argument

OpenStudy (compassionate):

I didn't comprehend that sentence at all.

OpenStudy (compassionate):

Wait...

OpenStudy (jdoe0001):

heheh

OpenStudy (compassionate):

Okay, so I'm going to work the problem by myself. Can you check it?

OpenStudy (jdoe0001):

\(\begin{array}{llll} f(\quad h(x)\quad )=\cfrac{12}{4({\color{brown}{ 3x+5}})+2}& h(\quad f(x)\quad )=3\left({\color{brown}{ \cfrac{12}{4x+2}}}\right)+5\\ \qquad \qquad \qquad \qquad \uparrow &\qquad \qquad \qquad \qquad \qquad \uparrow\\ \qquad \qquad \qquad \qquad f(x) &\qquad \qquad \qquad \qquad \qquad g(x) \end{array}\) see the function's insertion as the argument?

OpenStudy (jdoe0001):

woops... that should be h(x) anyhow.. not g(x) \(\begin{array}{llll} f(\quad h(x)\quad )=\cfrac{12}{4({\color{brown}{ 3x+5}})+2}& h(\quad f(x)\quad )=3\left({\color{brown}{ \cfrac{12}{4x+2}}}\right)+5\\ \qquad \qquad \qquad \qquad \uparrow &\qquad \qquad \qquad \qquad \qquad \uparrow\\ \qquad \qquad \qquad \qquad f(x) &\qquad \qquad \qquad \qquad \qquad h(x) \end{array}\)

OpenStudy (compassionate):

Why is f(x) being pointed to, when h(x) is 3x + 5, and why is there an arrow pointing as h(x) when that's f(x)...

OpenStudy (jdoe0001):

hmmm wait a sec. I got it backwards... shoot \(\begin{array}{llll} f(\quad h(x)\quad )=\cfrac{12}{4({\color{brown}{ 3x+5}})+2}& h(\quad f(x)\quad )=3\left({\color{brown}{ \cfrac{12}{4x+2}}}\right)+5\\ \qquad \qquad \qquad \qquad \uparrow &\qquad \qquad \qquad \qquad \qquad \uparrow\\ \qquad \qquad \qquad \qquad h(x) &\qquad \qquad \qquad \qquad \qquad f(x) \end{array}\) anyhow that's what f( h(x) ) and h( f(x) ) means

OpenStudy (compassionate):

I see - I see. So, you can kind of look at it. Since h(x) is on the inside, we can put it on the inside and surround it by f(x), yeah?

OpenStudy (jdoe0001):

right..... f ( h(x) ) means "replace any x's in the function with h(x) "

OpenStudy (compassionate):

Oh, I see (: Also, what were you saying about an inverse? That they can be equal?

OpenStudy (jdoe0001):

it just happen that this function only has one "x", thus only one replacement with h(x) if you had many "x", each would be replaced by h(x) since h(x) is the "argument"

OpenStudy (jdoe0001):

as far as I know, the only time you'd get the same number back from each case is only if the functions are inverse of each other

OpenStudy (jdoe0001):

\(\bf f(\quad h(x)\quad )=x\qquad h(\quad f(x)\quad )=x \\ \quad \\ only\ if\implies h(x)=f^{-1}(x)\quad and\quad f(x)=h^{-1}(x)\)

OpenStudy (compassionate):

Hmm. Thanks! I appreciate it!

OpenStudy (jdoe0001):

I don't think you're asked to get the inverse I gather just to explain whether they do yield the same value or not

OpenStudy (compassionate):

No, I wasn't. I appreciate your time!

OpenStudy (jdoe0001):

yw